OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

Nnesha (nnesha):

same set it equal to zero you will get quadratic equation \[\huge\rm Ax^2 + Bx+C =0\]

OpenStudy (anonymous):

so.. 7x^2-14x+2 = -5 then set it equal to zero by adding 5 to both sides. then factoring 7(x^2-2x+1) = 0 factor again 7(x-1)(x-1) = 0 so x = 1 with a multiplicity of 2?

Nnesha (nnesha):

oh nvm it's

Nnesha (nnesha):

yes right so now you can replace x by sin (sinx - 1)(sin x -1)= 0 so sin x =1 sin x = 1

Nnesha (nnesha):

and \[7 \cancel{=} 0\]